# Integrating Biological and Machine Intelligence: Attention Mechanisms in Brain-Computer Interfaces

### v1

@@ 3 Transformer-based Multi-Head Self-Attention Mechanisms in EEG
[42] 2022 Frequncy,Time Transformer seizure prediction CHB-MIT,Kaggle
[43] 2020 Time Transformer Sleep stage recognition MASS,Sleep-EDF
@@
[45] 2022 Time,Channel Transformer Visual discomfort induces private datasets
[46] 2021 Time,Channel Transformer Motor imagery BCI-IV 2A,BCI-IV 2B
[47] 2022 Time,Channel Transforemr Motor imagery PhysioNet
[48] 2023 Channel,Time Transformer emotion recognition DEEP,SEED,THU-EP
@@ 3.3 Practical Applications of Transformer Models in EEG Analysis
(1) Application in Temporal Dimension. In the temporal dimension, researchers have successfully leveraged CNN-Transformer models to extract valuable temporal information from EEG signals [56]. Enhancing the self-attention layer has enabled more effective capture of EEG patterns and detection of neural activity changes over time. For example, Li et al. [42] integrated CNN and Transformer architectures to adapt the self-attention mechanism to EEG-specific features. Unlike traditional approaches, their model segmented EEG signals into 30-second sequence samples and then generated time-frequency spectrograms by Short-Time Fourier Transform. These time-frequency spectrograms undergo preliminary feature extraction through multiple convolutional layers before entering an optimized Transformer module. Transformer module reduces the dimensions of the Query and Key matrices in the self-attention layer, lowering model complexity. Additionally, bias terms [19] were introduced to enhance positional information, further improving the model’s ability to handle complex temporal details in EEG signals. Similarly, Qu et al. adopted 30-second signal sequences as the model’s core units for sleep stage classification [43]. They implemented an encoder-decoder structure, beginning with a feature extraction layer combining convolution and pooling layers for initial preprocessing, followed by an encoder with two residual blocks to encode each signal segment. In the decoder, a Transformer-based self-attention mechanism was combined with fully connected layers to capture correlations across temporal sequence samples. The decoder processes these 30-sample subsequences by applying positional encoding to each sample before feeding it through the self-attention layer. The output then moves through two fully connected layers with normalization and dropout, culminating in sleep stage predictions via a softmax activation function. These studies demonstrate that Transformer-based self-attention mechanism can effectively extract dynamic changes in time-series EEG data, improving model efficiency and training speed through parallel computation.
@@
(2) Application in Spatial Dimension. Inspired by the Vision Transformer (ViT) model for image processing, researchers have developed innovative neural network models tailored for spatial feature analysis in EEG signals. For example, Guo et al. proposed a model called the Deep Convolutional and Transformer Network (DCoT) [44], focusing on the importance of each EEG channel in emotion recognition and visualizing these features. Based on the extracted differential entropy features, EEG signals were formed in a three-dimensional representation (Time $\times$ Channel $\times$ Frequency), which is similar to RGB representation in image. Inspired from ViT structure, DCoT was designed to analyze correlations between EEG electrodes, with each Transformer token representing a specific EEG electrode channel. The model applied positional encoding to the input tokens (i.e., EEG channels) and introduced an additional token dedicated to classification. After processing the encoded signals through the Transformer encoder module, a fully connected layer generated the final classification results. Furthermore, Zheng et al. proposed a Copula-Based Transformer model (CBT) [45], designed to learn spatial dependencies between EEG channels while optimizing classification performance. By reducing the size of the attention matrix, CBT lowered dependency on large datasets, improving computational efficiency. The CBT model excelled in an EEG-based visual discomfort assessment task, pioneering the revelation of temporal characteristics in EEG signals linked to visual discomfort. These studies underscore that Transformer-based self-attention mechanisms effectively capture the spatial dynamics of EEG signal sequences, enhancing spatial information processing and optimizing model training efficiency and accuracy.
@@
Researchers have leveraged the powerful capabilities of the Transformer self-attention mechanism to simultaneously capture both the temporal and spatial dimensions of EEG signals, enabling the extraction of highly discriminative features. For example, Song et al. [46] developed an EEG decoding model that primarily relies on the Transformer’s self-attention layers to enhance feature representations in both dimensions. In the spatial transformation component, the self-attention mechanism weighted each channel, emphasizing signals with higher relevance. Meanwhile, the temporal transformation component employed convolutional layers to encode temporal features, with channel compression and sample segmentation steps to reduce computational load. Finally, global average pooling and fully connected layers were combined to classify EEG signals effectively. Similarly, Du et al. proposed an EEG spatio-temporal Transformer network [47], which introduced Temporal Transformer Encoder (TTE) and Spatial Transformer Encoder (STE) modules to independently capture temporal and spatial features. In the TTE, temporal attention mechanisms calculated correlations among sampling points within each sample, extracting temporal features. Since channel correlations are often unique to individuals, the STE calculated spatial attention among channels and applies positional encoding to retain spatial location information. This allows the model to capture inter-channel relationships more accurately, improving its ability to identify individuals based on unique EEG patterns. Additionally, Si et al. [48] proposed integrating the Selective Kernel (SK) attention mechanism with the Transformer self-attention mechanism to identify and select the channels most relevant to the current task. This approach enables the model to filter out noise and irrelevant signals, simplifying the processing of complex temporal information. Once the SK attention mechanism isolated key channels, the Transformer encoder module deeply extracted temporal features from these selected channels. This approach’s strength lies in its ability to focus on the most informative parts of the data while preserving sensitivity to temporal dependencies, a critical factor for managing complex tasks in BCIs. These studies demonstrate that an incorporation of both temporal and spatial information can significantly enhance the accuracy and efficiency of EEG signal analysis.
@@ 4.1 Application of Attention Mechanisms in the Fusion of EEG Signals with Multimodal Data
Here, $\alpha_{1}$ and $\alpha_{2}$ represented the weight coefficients assigned to different modalities, satisfying $\alpha_{1}+\alpha_{2}=1$. The model’s performance was thoroughly evaluated by manually testing various combinations of $\alpha_{1}$ and $\alpha_{2}$. In the attention-based fusion method, although a weighted summation formula was still used to obtain the fused feature $O$, in solving for $\alpha_{1}$ and $\alpha_{2}$, the authors first initialized an attention layer, which was then optimized through the network. The final attention weights, $\alpha_{1}$ and $\alpha_{2}$, were derived from the softmax function applied to the attention layer’s output. This approach allows the model to adaptively adjust the attention weights, dynamically learning an optimal set of weight parameters.
On the other hand, Qiu et al. developed a correlation attention network that effectively integrated eye movement and EEG signals for fusion analysis. The unique aspect of the correlation attention network is that it not only computed attention weights for the two modalities but also adopted the canonical correlation analysis (CCA) method from statistics. This analysis method generated a cross-correlation coefficient matrix $C$ by calculating the cross-correlation between outputs of each recurrent unit in the bidirectional gated recurrent unit (GRU) layer. Utilizing this matrix along with the average features extracted by the bidirectional GRU, the fusion feature matrix $F$ was constructed to determine the attention weights. The computation followed a structured process. The attention score $u_{it}$ at each time step was derived as
@@
Additionally, Zhang et al. proposed a multimodal neural network model for integrating demographic data and EEG signals [63]. This model incorporated an attention mechanism to effectively combine the two data types, aiming to uncover complex relationships between EEG signals and demographic factors. Such integration is particularly valuable for applications like depression detection, where demographic information can provide crucial contextual insights. The model first used a one-dimensional CNN to process EEG signals, resulting in a feature matrix $Z\in\mathbb{R}^{C_{out}\times m}$, while demographic data were encoded as a feature vector $S\in\mathbb{R}^{d}$. The two data types were fused through an attention mechanism, with the calculation as follows:
@@
Recently, Transformer-based multi-head self-attention mechanisms have become a prevalent approach in multimodal data fusion, enabling efficient processing and integration of features from diverse modalities. In the model framework, each token received by the Transformer encoder corresponds to a specific modality, allowing the model to effectively capture and integrate information at the feature level. By leveraging complex inter-modal interactions, the self-attention mechanism uncovers intrinsic relationships and complementary information between modalities, dynamically refining feature representations. A well-known example is the MMASleepNet model [65], which not only incorporated the concept of the Squeeze-and-Excitation (SE) module from SENet [66] but also integrated the advantages of the Transformer encoder module. This combination enhances the model’s ability to effectively handle multimodal data by dynamically adjusting feature representations across different modalities, thereby improving overall performance.
@@
In EEG-related research, visual reconstruction from EEG signals is a particularly challenging yet rapidly evolving field. The objective is to decode the information embedded in EEG signals and use it to reconstruct corresponding visual stimuli. This process requires capturing the intricate relationships between EEG and visual modalities, often achieved through cross-attention mechanisms. By computing the relevance between EEG segments and visual data, EEG signals serve as conditioning inputs to guide the generation of visual stimuli. A well-known example is the DreamDiffusion model [71], which employed an EEG encoder trained on large-scale EEG data using a masking strategy to enhance feature extraction. The extracted EEG features then conditioned the Stable Diffusion model to generate images. Within Stable Diffusion, the cross-attention mechanism computes correlations between EEG-derived and image-derived features, enabling coherent information exchange between the two modalities. The mathematical formulation of cross-attention can be expressed as follows:
@@ 5 Conclusion and Future Works
This paper explores the integration of brain-computer interface (BCI) technology with attention models and their applications, highlighting recent advancements in the field. Traditional attention mechanisms, as a classical approach, are widely applied in EEG signal analysis. Transformer-based methods open up new avenues for processing and analyzing EEG signals, providing a more flexible framework. However, the effectiveness of attention mechanisms is often highly dependent on specific data types, limiting model generalization and posing challenges for cross-dataset transfer. These limitations emphasize the need for further empirical research to develop more generalizable and robust model architectures.
@@
In BCI applications, attention mechanisms not only expand research opportunities at the intersection of artificial intelligence and neuroscience but also enhance the technology’s potential and future prospects across various domains, including clinical applications, disability support, maritime safety, education, emotion recognition, and aerospace systems. Future research could primarily focus on the following aspects.
First, deep learning models, particularly those integrated with attention mechanisms, are expected to be applied in broader domains and tasks, such as cognitive neuroscience, clinical diagnostics, rehabilitation medicine, and affective computing. These applications can facilitate early detection of neurological disorders (e.g., Parkinson’s disease, epilepsy, and stroke recovery), improve assistive BCI systems for individuals with motor impairments, and enhance emotional state recognition for mental health monitoring. In education, BCI-driven attention-tracking and cognitive load assessment could support personalized learning experiences, benefiting students with learning disabilities such as ADHD. In maritime and aerospace industries, EEG-based fatigue monitoring systems could enhance situational awareness and reduce the risk of human errors in high-stress operational environments.
Second, more specialized model structures should be designed based on the characteristics of EEG signals. Currently, most model structures are adapted from fields like computer vision and natural language processing by converting EEG signals into image or text sequence data structures. However, for clinical neurotechnology, specialized architectures could improve real-time seizure prediction and brain-state monitoring for comatose patients. In disability support, adaptive BCI interfaces could enhance communication for individuals with locked-in syndrome or severe speech impairments. In maritime and aerospace neuroergonomics, models tailored to physiological constraints and environmental stressors could optimize pilot and ship crew performance, improving safety in high-risk settings.
Third, models employing self-attention layers based on Transformer architectures often have lower computational efficiency and slower convergence speeds. In future work, considering lightweight and energy-efficient model structures could be a promising direction. In disability assistance, real-time BCI-controlled prosthetics and exoskeletons require rapid response times to ensure seamless user interaction. In clinical applications, mobile and wearable BCI systems demand low-power, high-efficiency architectures for continuous neurophysiological monitoring. Similarly, in maritime and aerospace applications, where computational resources may be limited in remote or high-altitude environments, optimizing Transformer-based models for real-time decision-making could significantly enhance operational efficiency and safety.
In summary, the study of attention mechanisms will further advance the development of BCI technology. By integrating attention mechanisms with advanced technologies such as machine learning and deep learning, we can explore more sophisticated and domain-specific BCI systems. These innovations have significant implications for clinical rehabilitation, disability support, maritime safety, educational neuroscience, emotion-adaptive interfaces, and aerospace neuroergonomics, ultimately enabling safer, more adaptive, and more human-centric interactions between the brain and external systems.

### v2

@@ 3 Transformer-based Multi-Head Self-Attention Mechanisms in EEG
[67] 2020 Temporal Transformer Sleep stage recognition MASS, Sleep-EDF
[68] 2021 Temporal, Channel Transformer Motor imagery BCI IV-2A, BCI IV-2B
[69] 2022 Frequncy, Temporal Transformer Seizure prediction CHB-MIT, Kaggle datasets
@@
[71] 2022 Temporal, Channel Transformer Visual discomfort induces Private datasets
[72] 2022 Temporal, Channel Transformer Motor imagery PhysioNet
[73] 2024 Channel, Temporal Transformer Emotion recognition DEAP, SEED, THU-EP
[74] 2024 Channel, Frequncy Transformer Emotion recognition Private datasets, SEED, SEED-IV
[75] 2024 Temporal, Channel, Frequncy Transformer Emotion recognition, Motor imagery BCI IV-2A, BCI IV-2B, SEED
[76] 2025 Temporal, Channel Transformer Cognitive attention classification A publicly accessible EEG dataset
[77] 2025 Temporal Transformer EEG event recognition, epilepsy seizure detection TUEV, TUAB, IIIC Seizure, CHB-MIT
@@ 3.3 Practical Applications of Transformer Models in EEG Analysis
(1) Application in Temporal Dimension. In the temporal domain, researchers have increasingly adopted hybrid models such as CNN-Transformer and LSTM-Transformer to extract informative temporal representations from EEG signals [85]. Enhancements to the self-attention mechanism have significantly improved the models’ ability to capture evolving neural patterns and temporal dynamics. Qu et al. [67] proposed a residual-based attention model for sleep staging, treating 30-second EEG epochs as fundamental processing units. The model’s decoder utilized a multi-head self-attention mechanism, enriched with sinusoidal positional encoding, to retain the sequential structure of sleep transitions. By embedding residual blocks within the encoder for feature extraction and adopting parallelizable self-attention operations, the model effectively captured inter-epoch dependencies (e.g., sleep stage shifts) while mitigating the sequential processing limitations inherent in LSTM architectures. This design achieved a 12× acceleration in training speed, enabling efficient large-scale sleep data processing. Additionally, the model performed multi-scale frequency decomposition using Hilbert-transform-inspired pooling, and directly operated on raw EEG inputs in an end-to-end architecture, thus eliminating reliance on handcrafted preprocessing. Complementing this work, Li et al. [69] introduced the Transformer-guided CNN (TGCNN) for epilepsy prediction, featuring a novel squeezed multi-head self-attention (SMHSA) module. Unlike standard Transformer attention, SMHSA applied depth-wise convolutions to downsample the Key and Value matrices, reducing computational cost by a factor of four, while incorporating learnable bias terms to improve positional encoding. The model employed an alternating CNN-Transformer structure, in which CNN layers first extracted local time-frequency patterns from STFT-based spectrograms, followed by SMHSA modeling of long-range temporal dependencies. This allowed the model to capture both fine-grained waveform details (e.g., seizure spikes) and broader temporal trends (e.g., pre-ictal changes), achieving a robust balance between local feature resolution and global contextual awareness. While Qu et al. focused on optimizing training efficiency and temporal transition modeling for regular and stable sleep stages, Li et al. targeted computational scalability and feature complementarity to detect rare, transient epileptic events. However, each approach has its trade-offs: Qu’s reliance on fixed-length epochs may limit modeling of ultra-long temporal dependencies, whereas Li’s use of STFT preprocessing could result in loss of raw phase information. Together, these studies underscore the importance of task-specific adaptations of self-attention (such as residual-enhanced encoding in Qu’s model and compressed attention in Li’s architecture) in optimizing EEG temporal modeling. Future research directions may include the integration of adaptive temporal windows, multi-resolution temporal attention, or self-supervised pretraining to improve generalization across diverse EEG analysis tasks.
@@
(2) Application in Spatial Dimension. Inspired by the Vision Transformer (ViT) model for image processing, researchers have developed innovative neural network models tailored for spatial feature analysis in EEG signals. For example, Guo et al. proposed a model called the Deep Convolutional and Transformer Network (DCoT) [70], focusing on the importance of each EEG channel in emotion recognition and visualizing these features. Based on the extracted differential entropy features, EEG signals were formed in a three-dimensional representation (Time $\times$ Channel $\times$ Frequency), which is similar to RGB representation in image. Inspired from ViT structure, DCoT was designed to analyze correlations between EEG electrodes, with each Transformer token representing a specific EEG electrode channel. The model applied positional encoding to the input tokens (i.e., EEG channels) and introduced an additional token dedicated to classification. After processing the encoded signals through the Transformer encoder module, a fully connected layer generated the final classification results. Furthermore, Zheng et al. proposed a Copula-Based Transformer model (CBT) [71], designed to learn spatial dependencies between EEG channels while optimizing classification performance. By reducing the size of the attention matrix, CBT lowered dependency on large datasets, improving computational efficiency. The CBT model excelled in an EEG-based visual discomfort assessment task, pioneering the revelation of temporal characteristics in EEG signals linked to visual discomfort. These studies underscore that Transformer-based self-attention mechanisms effectively capture the spatial dynamics of EEG signal sequences, enhancing spatial information processing and optimizing model training efficiency and accuracy. These studies illustrate two complementary strategies for integrating spatial information into Transformer-based EEG models. DCoT adopts a data-driven self-attention paradigm, enabling the dynamic discovery of inter-channel correlations, and is particularly well-suited to applications demanding interpretability, such as emotion recognition with spatial brain mapping. In contrast, CBT leverages structured priors via copula functions to explicitly encode electrode positioning, which is advantageous in scenarios characterized by limited data availability and region-specific neural responses, such as visual discomfort detection. While both approaches validate the effectiveness of self-attention for spatial feature modeling, future research could explore hybrid frameworks that combine physiological priors (e.g., brain connectivity graphs, cortical maps) with adaptive attention mechanisms. Such integration may further enhance the generalization capability of EEG-based models across diverse cognitive and clinical tasks, especially when spatial specificity is critical.
@@
Researchers have leveraged the powerful capabilities of the Transformer self-attention mechanism to simultaneously capture both the temporal and spatial dimensions of EEG signals, enabling the extraction of highly discriminative features. For example, Song et al. [68] developed an EEG decoding model that primarily relies on the Transformer’s self-attention layers to enhance feature representations in both dimensions. In the spatial transformation component, the self-attention mechanism weighted each channel, emphasizing signals with higher relevance. Meanwhile, the temporal transformation component employed convolutional layers to encode temporal features, with channel compression and sample segmentation steps to reduce computational load. Finally, global average pooling and fully connected layers were combined to classify EEG signals effectively. Similarly, Du et al. proposed an EEG spatio-temporal Transformer network [72], which introduced Temporal Transformer Encoder (TTE) and Spatial Transformer Encoder (STE) modules to independently capture temporal and spatial features. In the TTE, temporal attention mechanisms calculated correlations among sampling points within each sample, extracting temporal features. Since channel correlations are often unique to individuals, the STE calculated spatial attention among channels and applies positional encoding to retain spatial location information. This allows the model to capture inter-channel relationships more accurately, improving its ability to identify individuals based on unique EEG patterns. Additionally, Si et al. [73] proposed integrating the Selective Kernel (SK) attention mechanism with the Transformer self-attention mechanism to identify and select the channels most relevant to the current task. This approach enables the model to filter out noise and irrelevant signals, simplifying the processing of complex temporal information. Once the SK attention mechanism isolated key channels, the Transformer encoder module deeply extracted temporal features from these selected channels. This approach’s strength lies in its ability to focus on the most informative parts of the data while preserving sensitivity to temporal dependencies, a critical factor for managing complex tasks in BCIs. These studies demonstrate that an incorporation of both temporal and spatial information can significantly enhance the accuracy and efficiency of EEG signal analysis. A critical factor for managing the variability and complexity of real-world BCI scenarios. Together, these studies illustrate the advantages of incorporating both temporal and spatial information within Transformer-based EEG models. Each approach presents task-specific strengths. Si et al. enable robust emotion recognition via hierarchical attention. Song et al. achieve efficient motor imagery decoding with lightweight slice-based self-attention. Du et al. enhance person identification using spatio-temporal encoders and positional encoding. These works highlight the growing trend of leveraging modular attention architectures to tailor EEG decoding strategies for diverse BCI applications.
@@ 4.1 Application of Attention Mechanisms in the Fusion of EEG Signals with Multimodal Data
Here, $\alpha_{1}$ and $\alpha_{2}$ represented the weight coefficients assigned to different modalities, satisfying $\alpha_{1}+\alpha_{2}=1$. The model’s performance was thoroughly evaluated by manually testing various combinations of $\alpha_{1}$ and $\alpha_{2}$. In the attention-based fusion method, although a weighted summation formula was still used to obtain the fused feature $O$, in solving for $\alpha_{1}$ and $\alpha_{2}$, the authors first initialized an attention layer, which was then optimized through the network. The final attention weights, $\alpha_{1}$ and $\alpha_{2}$, were derived from the softmax function applied to the attention layer’s output. This approach allows the model to adaptively adjust the attention weights, dynamically learning an optimal set of weight parameters. This work attempts to enhance multimodal fusion by incorporating adaptive weight coefficients. It systematically compares traditional manual weighting with an attention-based dynamic adjustment strategy, ultimately demonstrating improvements in both performance and robustness for multimodal emotion recognition. These findings further support the application potential of multimodal analysis in BCI systems. However, the model architecture was restricted to the fusion of only two modalities, and the employed attention mechanism was relatively simple in design. On the other hand, Qiu et al. designed a more complex attention structure. They developed a correlation attention network that effectively integrated eye movement and EEG signals for fusion analysis. The unique aspect of the correlation attention network is that it not only computed attention weights for the two modalities but also adopted the canonical correlation analysis (CCA) method from statistics. This analysis method generated a cross-correlation coefficient matrix $C$ by calculating the cross-correlation between outputs of each recurrent unit in the bidirectional gated recurrent unit (GRU) layer. Utilizing this matrix along with the average features extracted by the bidirectional GRU, the fusion feature matrix $F$ was constructed to determine the attention weights. The computation followed a structured process. The attention score $u_{it}$ at each time step was derived as
@@
Recent studies have pointed out that traditional multimodal fusion approaches tend to focus solely on integrating various types of physiological signal features. A novel and increasingly promising direction involves combining human statistical data (such as demographic information) with physiological signals from the brain. This integration provides a richer and more holistic understanding of brain activity and is expected to significantly advance the performance and applicability of BCI systems. Zhang et al. proposed a multimodal neural network model for integrating demographic data and EEG signals [92]. This model incorporated an attention mechanism to effectively combine the two data types, aiming to uncover complex relationships between EEG signals and demographic factors. Such integration is particularly valuable for applications like depression detection, where demographic information can provide crucial contextual insights. The model first used a one-dimensional CNN to process EEG signals, resulting in a feature matrix $Z\in\mathbb{R}^{C_{out}\times m}$, while demographic data were encoded as a feature vector $S\in\mathbb{R}^{d}$. The two data types were fused through an attention mechanism, with the calculation as follows:
@@
Recently, Transformer-based multi-head self-attention mechanisms have become a prevalent approach in multimodal data fusion, enabling efficient processing and integration of features from diverse modalities. In the model framework, each token received by the Transformer encoder corresponds to a specific modality, allowing the model to effectively capture and integrate information at the feature level. By leveraging complex inter-modal interactions, the self-attention mechanism uncovers intrinsic relationships and complementary information between modalities, dynamically refining feature representations. A well-known example is the MMASleepNet model [94], which not only incorporated the concept of the Squeeze-and-Excitation (SE) module from SENet [95] but also integrated the advantages of the Transformer encoder module. To enhance the utilization of multimodal collaborative information, researchers typically employ a two-stage attention mechanism: first, applying attention modules to extract modality-specific features from individual modalities, followed by integrating cross-modal information through another layer of attention during fusion. This strategy, which combines hierarchical attention mechanisms, has demonstrated remarkable effectiveness in BCI tasks. For instance, He et al. proposed the EEG-EMG FAConformer framework[96], which leverages self-attention modules to hierarchically extract critical features from EEG and EMG(electromyography) signals. By further incorporating cross-modal attention for fusion, this approach achieves robust performance and enhanced classification accuracy, as evidenced by its superior stability and generalization capabilities in experimental evaluations. Another interesting research work by Wang et al. proposed a cross-modal fusion model to combine EEG signals with textual data for enhancing emotion detection in English writing[97]. The model utilizes the self-attention mechanism of the Transformer architecture to dynamically fuse EEG signals and textual features, addressing the limitations of traditional methods in leveraging the complementary information between textual and physiological signals.
@@
In EEG-related research, visual reconstruction from EEG signals is a particularly challenging yet rapidly evolving field. The objective is to decode the information embedded in EEG signals and use it to reconstruct corresponding visual stimuli. This process requires capturing the intricate relationships between EEG and visual modalities, often achieved through cross-attention mechanisms. By computing the relevance between EEG segments and visual data, EEG signals serve as conditioning inputs to guide the generation of visual stimuli. A well-known example is the DreamDiffusion model [102], which employed an EEG encoder trained on large-scale EEG data using a masking strategy to enhance feature extraction. The extracted EEG features then conditioned the Stable Diffusion model to generate images. Within Stable Diffusion, the cross-attention mechanism computes correlations between EEG-derived and image-derived features, enabling coherent information exchange between the two modalities. The cross-attention mechanism employed in Stable Diffusion model has been widely adopted in EEG-to-image guided reconstruction tasks, such as in works like[103, 104]. The mathematical formulation of cross-attention can be expressed as follows:
@@
4.3 Comparison of Transformer-Based Cross-Attention and Traditional Attention in EEG Multimodal Applications
Traditional attention mechanisms and Transformer-based cross-attention have both been applied in EEG multimodal fusion tasks. In recent years, Transformer-based approaches have increasingly become the mainstream. Traditional attention retains limited advantages in computational efficiency and parameter compactness, often using fewer parameters, which suits real-time BCI and low-resource clinical applications[105]. However, Transformer-based models demonstrate superior overall performance. For example, the Dual-Branch Transformer with Cross-Attention (DTCA) achieves state-of-the-art results on SEED and SEED-IV datasets by encoding EEG and eye movement data with sophisticated query-key-value relationships[106]. In more complex scenarios involving a greater number of modalities, Transformer-based cross-attention mechanisms also demonstrate strong performance. For example, Ying et al. developed a method that integrates EEG, audio, and video, advancing the incorporation of EEG into more complex multimodal systems[107]. Notably, researchers are actively exploring lightweight Transformer architectures. In the future, significant improvements in the computational efficiency of Transformer-based cross-attention mechanisms are expected to further reduce the reliance on traditional attention in multimodal EEG systems.
@@ 5 Conclusion and Future Works
This paper presents a systematic and comprehensive review of recent advancements in attention mechanisms integrated into EEG-based BCI systems. We structure the review around three conceptually related categories of attention mechanisms: traditional attention methods, Transformer-based multi-head self-attention architectures, and multimodal attention fusion strategies. First, we introduce traditional attention mechanisms, including channel-specific, temporal-specific, and frequency-specific modules. These approaches enhance EEG signal decoding by adaptively emphasizing salient features, which could improve model accuracy and robustness against noise and inter-subject variability. We then explore Transformer-based architectures that utilize multi-head self-attention mechanisms to capture complex temporal and spatial dependencies within EEG data. These methods enable more expressive and flexible feature representations, addressing limitations of earlier approaches. We then review multimodal attention fusion strategies, which incorporate EEG signals with additional physiological and sensory modalities. This integration enhances system interpretability and adaptability, supporting more effective performance in complex and dynamic environments. Together, these three categories form a coherent taxonomy that reflects the methodological evolution of attention-based techniques in neural decoding and offers a comparative understanding of their respective advantages, limitations, and suitable application contexts. Through this structured overview, we aim to provide researchers with a practical reference for selecting attention mechanisms that best address specific challenges in BCI design.
Despite the progress outlined in this review, several important challenges remain. Current attention mechanisms often suffer from limited generalizability, as their performance tends to be constrained by dataset-specific and subject-specific characteristics. In addition, the high computational complexity of models such as Transformers can hinder their application in real-time and resource-limited scenarios. To address these issues, future research should focus on the development of lightweight, energy-efficient, and EEG-specific attention models that are better aligned with the physiological properties of neural signals. Advancing in this direction will help unlock the broader potential of attention-based BCIs, enabling their practical use across a range of domains including clinical diagnosis, neurorehabilitation, education, and human–machine interaction.
To establish a stronger foundation for future advancements, it is essential to examine the current landscape of attention-based BCI applications across various domains. In clinical neuroscience, attention mechanisms have significantly enhanced diagnostic accuracy in EEG-based emotion recognition and the detection of neurological disorders such as epilepsy and Parkinson’s disease. However, these clinical systems continue to face challenges, particularly with inter-subject variability and the need for extensive calibration, which hinder widespread clinical deployment. In motor rehabilitation, attention mechanisms optimized for specific EEG channels and temporal dynamics have improved the decoding of motor imagery. Nonetheless, achieving robust, low-latency performance suitable for real-time application in diverse operational environments remains a key challenge. In educational technology, attention-based BCIs have made substantial progress in enabling personalized learning, particularly for students with learning disabilities. However, the scalability of such systems and their integration into formal educational settings still require significant development. In safety-critical industries, such as aviation and maritime operations, EEG-driven attention monitoring has shown promise in managing fatigue and cognitive workload. Despite this potential, ensuring system reliability and generalizability under dynamic, high-stress conditions remains a critical concern.
Building upon these practical experiences and the challenges identified above, future research in attention-based deep learning for BCIs is well-positioned to address current limitations and expand into broader domains with increased precision and robustness. Specifically, attention-enhanced models have the potential to reshape several application areas by leveraging the temporal-spatial sensitivity of EEG signals and the interpretability of attention mechanisms. In clinical neuroscience, attention-guided BCI frameworks can be utilized to develop early diagnostic tools for conditions such as epilepsy, Parkinson’s disease, and Alzheimer’s disease. By dynamically attending to frequency bands and electrode regions associated with preictal states, attention-based models can facilitate real-time seizure prediction with fewer false alarms. In assistive technologies, attention mechanisms can enhance brain-controlled exoskeletons and spelling interfaces for individuals with severe motor impairments. Through adaptively focusing on high-informative EEG segments, such as event-related desynchronization (ERD) windows, the system could improve command accuracy under non-stationary conditions. Coupling these models with transfer learning and few-shot adaptation methods may reduce calibration time and personalize device control for locked-in syndrome patients. In mental health monitoring, attention-based emotion recognition systems can be fine-tuned to identify neural signatures of chronic stress, anxiety, or depression. Temporal attention can be applied to isolate emotionally salient EEG segments during resting-state or affective stimulation, facilitating early intervention via wearable neurofeedback devices or closed-loop neuromodulation. In educational neuroscience, future BCI systems can employ attention modules to track cognitive workload in real time, enabling adaptive learning platforms that adjust content difficulty or presentation speed based on attention levels and fatigue. These systems could be particularly beneficial for students with ADHD, where continuous monitoring of frontoparietal EEG activity might allow for real-time feedback or stimulus modulation to improve engagement and learning outcomes. In maritime and aerospace environments, attention-informed BCIs can support neuroergonomic monitoring by detecting shifts in vigilance, workload, or mental fatigue during prolonged or high-stakes operations. By focusing on alpha and theta band dynamics in frontal and parietal regions, these systems can be trained to issue cognitive alerts or task reallocations in response to declining neural performance. Furthermore, embedding these models in low-power wearable headsets with real-time edge inference capabilities is essential for deployment in dynamic, mobility-constrained settings.
Furthermore, future research on attention-based BCI systems should prioritize the development of specialized model architectures that are explicitly aligned with the unique physiological and temporal characteristics of EEG signals. At present, many BCI models are adapted from computer vision and natural language processing, relying on transformations that convert EEG signals into images or sequences. While these approaches have shown promise, they often fail to fully capture the dynamic and non-stationary nature of neural activity. Designing models specifically tailored to EEG data could lead to substantial improvements in clinical neurotechnologies, such as real-time seizure prediction and continuous brain-state monitoring. In addition, it is critical to address the computational inefficiencies and convergence challenges commonly associated with Transformer-based attention mechanisms. These limitations are particularly relevant for mobile neurodiagnostic tools, wearable assistive technologies, and real-time neuroergonomic monitoring systems, where energy efficiency, low latency, and portability are essential. Developing lightweight and optimized attention models will be vital not only for enhancing user experience but also for ensuring reliable real-time inference in resource-constrained environments.